A 240-V clothes dryer draws a current of 15 A. How much energy, in kilowatt-hours and in joules, does it use in 45 minutes of operation?


1
Expert's answer
2021-05-25T16:29:10-0400

Gives

Voltage (v)=240V

Current (i)=15A

Time=45min=0.75h

We know that

"P=\frac{E}{t}"

"E={P}\times{t}"

Where P=power

E=energy

t=time

When

P"=IV"

Put value

"P=15\times240=3600W"

"E=3600\times0.75Wh"

"E=2700Wh"

Energy found kWh

"E=\frac{2700}{1000}=2.7KWh"

1KWh="3.6\times10^6J"

E="2.7\times3.6\times10^{6}J"

"E=9.72\times20^6J"